OpenStudy (anonymous):

x+y=5 and xy=-1find the value of x^2 +y^2

OpenStudy (anonymous):

\[x^2+y^2=x^2+2xy+y^2-2xy=(x+y)^2-2xy\] Now everything you need to do is substitute the values. Hope this help.